OpenStudy (anonymous):

The total revenues R(in thousands of dollars) earned from manufacturing and selling hand-held video games is given by R(p)=-25p^2+1200p, where p is the price per unit(in dollars). What is the unit price that will yield maximum revenue?

OpenStudy (amistre64):

aha!! derivatives in action.... maybe :)

OpenStudy (amistre64):

Anwar can get this right?

OpenStudy (amistre64):

hmm....or I can lol

OpenStudy (anonymous):

I see my name :P

OpenStudy (amistre64):

R' = -50p +1200 = 0 p = 1200/50 = 120/5 = 24 when the price is at 24 you get a maximum yeild in revenue.... i need a double check on that ;)

OpenStudy (anonymous):

I am sure that you can

OpenStudy (amistre64):

we could also have gone with the quad formulas -b/2a to find the center of the hump... -1200/-50 = same results

OpenStudy (anonymous):

thanks guys..

OpenStudy (amistre64):

youre welcome :)